Each line of the following table gives data for a process of a closed system. Each entry has the same energy unit. Determine the missing entries. Write down the procedure for each part.


ProcessQWE1E2Δ E
A+50?-20?+70
B?+20?+50+30
C?-60+40+60?
D-40?+50?0
E+50+150?-80?
Answer #1

Process.A:

Apply energy balance,

$$ \begin{aligned} &Q=W+\Delta E \\ &50=W+70 \\ &W=-20 \end{aligned} $$

Find the value of \(E_{2}\).

$$ \begin{aligned} &\Delta E=E_{2}-E_{1} \\ &70=E_{2}-(-20) \\ &E_{1}=50 \end{aligned} $$

Process. B:

Apply energy balance,

$$ \begin{aligned} &Q=W+\Delta E \\ &Q=20+30 \\ &Q=50 \end{aligned} $$

Find the value of \(E_{1}\)

$$ \begin{aligned} &\Delta E=E_{2}-E_{1} \\ &30=50-E_{1} \\ &E_{1}=20 \end{aligned} $$

Process C

Find the value of \(\Delta E\).

$$ \begin{aligned} \Delta E &=E_{2}-E_{1} \\ &=60-40 \\ \Delta E &=20 \end{aligned} $$

Apply energy equation.

$$ \begin{aligned} &Q=W+\Delta E \\ &Q=-60+20 \\ &Q=-40 \end{aligned} $$

Process D:

Find the value of \(E_{2}\).

$$ \begin{aligned} &\Delta E=E_{2}-E_{1} \\ &0=E_{2}-50 \\ &E_{2}=50 \end{aligned} $$

A pply energy equation.

$$ \begin{aligned} &Q=W+\Delta E \\ &-40=W+0 \\ &W=-40 \end{aligned} $$

Process E:

Apply cnergy equation.

$$ \begin{aligned} &Q=W+\Delta E \\ &50=150+\Delta E \\ &\Delta E=-100 \end{aligned} $$

Find the value of \(E_{1}\).

$$ \begin{aligned} &\Delta E=E_{2}-E \\ &-100=-80-E_{1} \\ &E_{1}=20 \end{aligned} $$